OpenStudy (anonymous):

4 over the square root of 21

OpenStudy (jdoe0001):

ok...... so...?

OpenStudy (anonymous):

Can you write the equation?

OpenStudy (jdoe0001):

\(\bf \cfrac{4}{\sqrt{21}}\) I'd think, but that means nothing anymore than just saying.... 27

OpenStudy (anonymous):

simplify the radical expression by rationalizing the denominator 4 over the square root of 21

OpenStudy (anonymous):

it is the same thing \[4 (21)^{-1/2}\]

OpenStudy (anonymous):

But first do the exponent then multiply by 4

OpenStudy (anonymous):

just simplify it please

OpenStudy (jdoe0001):

ohhh.. ahemm ok. well, simplifying a rational with a radical in the denominator means, getting rid of the root at the bottom, and you'd do that by multiplying top and bottom by the denominator, that is \(\bf \cfrac{4}{\sqrt{21}}\cdot \cfrac{\sqrt{21}}{\sqrt{21}}\implies \cfrac{4\sqrt{21}}{(\sqrt{21})^2}\implies \cfrac{4\sqrt{21}}{21}\)

OpenStudy (anonymous):

and then what do u do after that

OpenStudy (jdoe0001):

well, you can't divide top and bottom because they have no common factors, so that's as much as you'd simplify it

OpenStudy (anonymous):

how bout this one

OpenStudy (anonymous):

8 over radical 6 - radical 3

OpenStudy (jdoe0001):

do you know what a "conjugate" is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but can you please just solve it for me

OpenStudy (anonymous):

please

OpenStudy (jdoe0001):

http://www.mathsisfun.com/algebra/conjugate.html <---- conjugate gimme a sec, let us use the conjugate of the denominator

OpenStudy (anonymous):

i just really need the answer asap

OpenStudy (jdoe0001):

\(\bf{\color{blue}{ (a-b)(a+b) = a^2-b^2}} \\ \quad \\ \quad \\ \cfrac{8}{\sqrt{6}-\sqrt{3}}\cdot \cfrac{\sqrt{6}+\sqrt{3}}{\sqrt{6}+\sqrt{3}}\implies \cfrac{8(\sqrt{6}+\sqrt{3})}{(\sqrt{6}-\sqrt{3})(\sqrt{6}+\sqrt{3})} \\ \quad \\ \cfrac{8(\sqrt{6}+\sqrt{3})}{(\sqrt{6})^2-(\sqrt{3})^2}\)

OpenStudy (anonymous):

final answer?

OpenStudy (jdoe0001):

simplify the rational, as you can see the denominator comes out of the radical

OpenStudy (anonymous):

ive been trying to do that but cant, please please give me the denominator

OpenStudy (anonymous):

plssss

OpenStudy (anonymous):

is the denominator 27?????

OpenStudy (jdoe0001):

recall that \(\bf \Large \sqrt{x}\implies \sqrt[{\color{red}{ 2}}]{x^{\color{red}{ 2}}}\implies x\)

OpenStudy (anonymous):

??? is the denominator 27?

OpenStudy (jdoe0001):

recall your radicals -> \(\bf \large (\sqrt{6})^2\implies (\sqrt[2]{6})^2\implies \sqrt[2]{6^2}\)

OpenStudy (anonymous):

so whats the final answer, just give me that pleaseee

OpenStudy (anonymous):

PLEASEEEEE

OpenStudy (jdoe0001):

that removes the neccessity for the exercise

OpenStudy (anonymous):

please, im begging you, just give me the answer pleaseeeeeee

OpenStudy (anonymous):

just tell me the final answer pleaseeeeeee

OpenStudy (anonymous):

ive been stuck for hours

OpenStudy (anonymous):

????? please just give me the final answer

OpenStudy (anonymous):

.....